rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#12045 — FS#16364 — CHI <> PAL

Attached to Project— Network
Incident
Backbone US
CLOSED
100%
We have 4x10G and 6x10G down between our POPs of Chicago and Palo-Alto.

We are redirecting part of the traffic to relieve the links LAX <> SNJ and will contact our supplier.
Date:  Monday, 01 February 2016, 18:12PM
Reason for closing:  Done
Comment by OVH - Monday, 01 February 2016, 18:09PM

01/28/2016 20:36:37 GMT - The Fiber Splice crew has arrived at the cut site. The fiber will need to be excavated approximately 20 feet back from the man hole to create enough slack for repairs. There remains no ETR to provide at this time.

01/28/2016 19:47:51 GMT - Field Services have identified the fiber cut in Leavenworth, KS and a fiber splice crew is en route. Fiber prep work has begun but there is no ETR available at this time.

01/28/2016 18:58:19 GMT - The Transport NOC has reported a possible fiber cut between Kansas City, MO and Lancaster, KS is impacting unprotected transport services. Field Services have been engaged to investigate with no ETR to provide at this time. Please be advised that updates for this event will be relayed on an hourly basis unless otherwise noted.


Comment by OVH - Monday, 01 February 2016, 18:09PM

01/28/2016 21:02:34 GMT - The NOC has advised that the fiber excavation work, to provide slack sufficient for repairs, is underway. Splice work is currently expected to begin at 22:00 GMT.

01/28/2016 21:51:52 GMT - Field Services reports that fiber prep work is ongoing and splicing is still expected to begin at approximately 22:00 GMT. A second splicer is being dispatched to assist.


Comment by OVH - Monday, 01 February 2016, 18:10PM

01/28/2016 21:02:34 GMT - The NOC has advised that the fiber excavation work, to provide slack sufficient for repairs, is underway. Splice work is currently expected to begin at 22:00 GMT.

01/28/2016 21:51:52 GMT - Field Services reports that fiber prep work is ongoing and splicing is still expected to begin at approximately 22:00 GMT. A second splicer is being dispatched to assist.


Comment by OVH - Monday, 01 February 2016, 18:10PM

28-Jan-2016 23:30:55 GMT - Field Services has confirmed that splice work is ongoing and that two splice crews are working to complete repairs as quickly as possible. Due to the extent of the cut and necessary splice work, repair efforts may take several hours to complete, with no ETR available. Individual services may restore as the splice work progresses. Please be advised that further updates will be provided as new information becomes available.


Comment by OVH - Monday, 01 February 2016, 18:12PM

All of our links are up. We are reactivating the traffic on the POPs of Palo-Alto and San Jose.


Comment by OVH - Monday, 01 February 2016, 18:12PM

Traffic has been reactivated.